# Modified Debye-Hückel-Onsager Theory for Electrical Conductivity in Aqueous Electrolyte Solutions. Nonlocal Electrostatics Approach

### v1

@@ Abstract
The paper presents a nonlocal mean field theory of electrolyte solutions, extending the classical Debye-Hückel-Onsager theory to provide a detailed description of the electrical conductivity in strong electrolyte solutions. The theory systematically incorporates the effects of ions specificity, such as steric interactions, hydration of ions, and their spatial charge distributions, into the mean-field framework. This allows for calculation of ion mobility and electrical conductivity, while accounting for relaxation and hydrodynamic phenomena. At low concentrations, the model reproduces the well-known Kohlrausch’s limiting law. Using the exponential (Slater-type) charge distribution function for solvated ions, we demonstrate that experimental data on the electrical conductivity of aqueous 1:1, 2:1, and 3:1 electrolyte solutions can be approximated over a broad concentration range by adjusting a single free parameter representing the spatial scale of the nonlocal ion charge distribution. Using the fitted value of this parameter at 298.15 K, we obtain good agreement with available experimental data when calculating the electrical conductivity across different temperatures. We also analyze the effects of temperature and electrolyte concentration on the relaxation and electrophoretic contributions to total electrical conductivity, explaining the underlying physical mechanisms responsible for the observed behavior.
@@ I Introduction
Electrolyte solutions [1] are essential for maintaining proper function and balance within living organisms, as they contain ions crucial for cellular processes such as nerve signaling [2, 3], muscle contraction [4, 5], and fluid balance [6]. These solutions can be found in natural sources like fruits and vegetables, as well as in sports drinks and oral rehydration solutions. Beyond biology and the food industry, electrolytes are critical components in technologies with significant societal and environmental impacts. In battery technology, electrolytes facilitate the movement of ions between electrodes, allowing energy storage and release [7, 8, 9, 10, 11]. With the growing global demand for fresh water, the role of electrolytes in desalination processes [12, 13] is becoming increasingly important. Electrolytes play a key role in technologies such as reverse osmosis and electrodialysis, facilitating the removal of contaminants from water and the production of clean drinking water. Thus, predicting electrical conductivity in solutions containing strong electrolytes poses significant challenges in physical chemistry and chemical physics, with implications for modern biology, medicine, food industry, agriculture, electronics, water purification, and environmental monitoring technologies.
@@
Debye, Hückel, Onsager, Fuoss, and Falkenhagen [14, 15, 16, 17, 18, 19, 20] laid the foundation for understanding the statistical theory of electrical conductivity in electrolytes. More recently, renewed interest and progress in this field have arisen from advancements in the statistical theory of liquids and liquid-phase solutions since the publication of these seminal papers. Approaches based on integral equations for distribution functions [21, 22, 23], stochastic density functional theory (SDFT) [24, 25, 26, 21], and self-consistent field theory based on the Poisson-Boltzmann equation [27, 28] have been developed and applied with varying degrees of success to model the electrical conductivity of electrolyte solutions across a broad range of concentrations and temperatures.
@@
Let us review the current state of the art in this field. The paper [25] by Avni et al. examines the conductivity of ionic solutions using SDFT combined with a modified Coulomb interaction that takes into account the hard-core repulsion between ions. The authors claim that the modified potential suppresses short-range electrostatic interactions present in the original Debye-Hückel-Onsager theory, leading to improved conductivity results. The conductivity values obtained from the model agree well with the experimental data up to a concentration of 3 M. In addition, the authors provide a compact expression and a simple analytical approximation for the conductivity. In a recent paper by Avni et al. [26], the authors extended their study to include multivalent ions and high electric fields, where a deviation from Ohm’s law, known as the Wien effect, occurs. In article [21], the authors explore recent approximations for computing the conductivity and self-diffusion coefficient of electrolytes, with a focus on the specifics of truncating Coulomb interactions at short distances. They compare these new approaches to traditional methods, such as the mean-spherical approximation and mode-coupling theory. The authors also discuss the importance of considering hydrodynamic effects in SDFT and demonstrate how the choice of modified Coulomb interactions can affect the properties of electrolytes. In general, the authors highlight the need for continued development and refinement of SDFT approaches for studying electrolytes and emphasize the valuable insights that can be gained by comparing these methods with other theoretical frameworks. In study [22], the authors developed a new electrical conductivity model for non-associated electrolyte solutions based on the Debye-Hückel-Onsager theory. This model assumes a single cation and a single anion within a continuum medium of the solvent(s), utilizing their crystallographic ionic radii. The authors compared the predictions of their model with experimental measurements of various binary aqueous solutions over a wide temperature range, from 273.15 to 373.15 K, and demonstrated good agreement with the experimental data. In paper [27], the authors provided a comprehensive examination of the unique behavior of electrolyte solutions conductivity when confined to micro- and nanochannels, as compared to bulk phases. They focuses on the reduction of ion mobility with increasing salt concentration, which impacts the conductivity of monovalent solutions as they transition from bulk to confinement in a narrow slit. The authors systematically treated electrophoresis of ions, deriving equations for zeta potentials and ion mobilities, and subsequently developed a simplified expression for bulk conductivity. In the paper [28], the same authors demonstrated that the derived expression accurately describes the conductivity of univalent salt solutions. Note that this model accounts only for a mean-field approximation of electrostatic interactions and utilizes an adapted classical approach to calculate colloid electrophoresis, as described in [29], without considering the relaxation effect of salt ions.
@@
Despite several attempts to theoretically describe electrical conductivity in aqueous solutions of strong electrolytes, some important issues remain unclear from the perspective of chemical physics and physical chemistry. Firstly, to the best of our knowledge, the nonlocal nature of the ion charge distribution has not been systematically explored in relation to the electrical conductivity of electrolyte solutions. More specifically, according to quantum theory, the charge of solvated ions (with their surrounding solvent molecules in solvation shell) is distributed over a region rather than being concentrated at a single point or spread over a thin shell. In other words, from the perspective of electrostatics, it should behave more like a smooth cloud than a point-like charge or a uniformly charged spherical shell. Based on general considerations, this effect is expected to become significant at sufficiently high solution concentrations, where the interactions between ions can no longer be accurately described as interactions between the point-like charges. It should be noted that recent studies have effectively attempted to account for charge nonlocality by modifying the Coulomb potential at short distances [26, 25, 21]. However, these modifications have not yet been explicitly related to the nature of the charge distribution nonlocality. Note that renormalization of the Coulomb potential at short distances has been systematically considered in plasma physics taking into account quantum effects related to the Heisenberg uncertainty relation [30]. Quantum mechanical calculations of the effective electrostatic interactions between ions in electrolyte solutions present a greater challenge than those in plasma physics, primary due to the presence of multiple electrons in solvated ions. To address this, computationally expensive ab initio molecular dynamics simulations could be used. However, for practical calculations of electrolyte solutions, it is also possible to use phenomenological models of the charge distribution of solvated ions adopted in phenomenological nonlocal electrostatics approach [31, 32, 33, 34] or within the microscopic statistical field theory [35, 36, 37, 38, 39]. These approaches lead to a truncation of the Coulomb potential at short distances. Secondly, the effect of ion hydration on the phenomenon of ion transport in a solution under an external electric field is not fully understood. Quantitatively, this effect should be analyzed by considering the effective size of the ions (ion hydration radii), taking into account the steric interactions between the ions and their effect on the dielectric constant of the solution (dielectric decrements [40, 41]). For this purpose, steric interactions between ions at close distances could be modeled as interactions of hard spheres with diameter equal to twice the hydration (solvation) radius. It is important to note that existing theoretical models [21, 26, 25] take into account steric effects using crystallographic radii of ions, rather than solvation radii.
@@
This study proposes a modified version of the Debye-Hückel-Onsager (DHO) theory that takes into account ion solvation, steric interactions between ions, and the non-local distribution of charge on solvated ions. We will utilise this theory to derive analytical expressions for the electrophoretic and relaxation contributions to ion mobility, as well as to describe the electrical conductivity of electrolyte solutions containing ions of different valencies across a wide range of concentrations and temperatures.
@@ II Statement of problem
With a very diluted solution, the mobility of individual ions approaches a constant limit given by the equation:
@@
For spherical ions, the viscous friction coefficient can be expressed as $\gamma_{a}=6\pi\eta R_{a}$, where $R_{a}$ represents the hydrodynamic (Stokes) radius of the $a$-th ion, and $\eta$ denotes the shear viscosity of the solution. In solutions at finite concentrations, each ion is surrounded by a screening charge cloud (atmosphere), which alters its mobility due to two distinct effects. The motion of ions in an external field causes a distortion in the charge distribution within the cloud, leading to an additional ’relaxation’ field acting on the ion. Furthermore, the movement of the ionic atmosphere induces motion in the solvent, resulting in the ion being "carried off". The impact of the first type of contribution on mobility is described as relaxation[42, 43, 20], while the second type is referred to as electrophoretic [42, 43, 27]. In the following, we will present an approach to calculating both contributions to the electrical conductivity of strong electrolyte solutions.
@@ III Relaxation contribution
III.1 General theory
Let us begin by calculating the relaxation contribution to the ion mobilities. We can define the two-point ion concentrations, denoted as $c_{ab}(\mathbb{r}_{a},\mathbb{r}_{b})$. The value $c_{ab}(\mathbb{r}_{a},\mathbb{r}_{b})d\mathbb{r}_{a}$ represents the number of ions of type $a$ near point $\mathbb{r}_{a}$ when a test ion of type $b$ is fixed at point $\mathbb{r}_{b}$. It is important to note that the ions can be of the same or different types, so we have the following symmetry relation:
@@ III.1 General theory
In the configuration space of two ions, $a$ and $b$, in a steady state, which we will only consider in this work, the following continuity equation holds [42, 19, 20]
$\nabla_{a}\mathbb{j}_{a}+\nabla_{b}\mathbb{j}_{b}=0,$ (5)
@@
[…]
is the smeared electrostatic potential at point $\mathbb{r}_{a}$ when an ion with total charge $q_{b}$ is located at point $\mathbb{r}_{b}$. The first term in equations (6) and (7) describes the diffusion of ions $a$ even in the absence of an external field. The second term describes the particle flux density due to the action of the electrostatic forces from the external field and the ’molecular’ electric field $-\nabla_{a}\bar{\psi}_{b}$ created at point $\mathbf{r}_{a}$ by the distorted ionic cloud, assuming that ion $b$ is located at point $\mathbf{r}_{b}$. The local electrostatic potential, $\psi_{b}=\psi_{b}(\mathbb{r}_{a},\mathbb{r}_{b})$, satisfies the standard Poisson equation
@@
[…]
@@
[…]
@@
where $J_{al}=\partial\mu_{a}/\partial c_{l}$, with $\mu_{a}=\mu_{a}(\{c_{l}\})$ being the equilibrium bulk chemical potentials of species. Note also that due to translation invariance all functions should depend only on the vector $\mathbb{r}=\mathbb{r}_{a}-\mathbb{r}_{b}$, so that
@@
The system of coupled equations (15) and (16) together form the basis of the modified DHO theory within a nonlocal electrostatics approach, which we will use in what follows to calculate the relaxation contribution to ion mobilities and electrical conductivity.
In accordance with the Onsager approach[20, 42], we find the solutions in the form
@@
where we introduced the reciprocal matrix, $\hat{J}_{ab}=(J^{-1})_{ab}$. The equilibrium electrostatic potential can be obtained from equation
@@
$\nabla^{2}G(\mathbb{r})-\mathcal{K}G(\mathbb{r})=-\frac{1}{\varepsilon\varepsilon_{0}}\varrho(\mathbb{r}),$ (26)
@@
For point-like charges for which $\varrho_{\mathbb{k}}=1$ we obtain
@@
where $r=|\mathbb{r}|$. Thus, $\kappa^{-1}$ in this case can be interpreted as the screening length [47]. Its dependence on concentrations and temperature is determined by the dependence of the bulk chemical potentials on concentrations, i.e. $\mu_{a}=\mu_{a}(\{c_{l}\})$ which we will specify below. The equilibrium perturbation concentrations, thereby, are
@@
The steady-state perturbations $\bar{c}_{ab}^{(1)}(\mathbb{r})$ and $\psi_{b}^{(1)}(\mathbb{r})$ which are the scalar values have the structure $(\mathbb{E}\cdot\mathbb{r})f(r)$, where $f(r)$ is the function only of the absolute value of $\mathbb{r}$. Therefore, $\bar{c}_{ab}^{(1)}(\mathbb{r})$ and $\psi_{b}^{(1)}(\mathbb{r})$ are odd functions. Bearing in mind condition (4), we arrive at the following property.
@@
$\nabla^{2}\psi_{b}^{(1)}(\mathbb{r})=-\frac{1}{\varepsilon\varepsilon_{0}}\sum\limits_{l}q_{l}\varrho\bar{c}_{lb}^{(1)}(\mathbb{r}),$ (33)
III.2 Application to one-component electrolyte solutions
Now, as an application of the proposed theory, let us consider only the case of one-component electrolyte solutions for which indices $a,b=1,2$, so that $q_{1}=z_{1}q$, $q_{2}=-z_{2}q$, $z_{a}=|q_{a}|/q$ is the charge number of ion $a$, $q=e$ is the elementary charge. In this case, only one variable $\bar{c}_{12}^{(1)}(\mathbb{r})\neq 0$. The basic equations (32) and (33) take the form
where we have taken into account that $\hat{J}_{11}=J_{22}/|J|$, $\hat{J}_{22}=J_{11}/|J|$, $\hat{J}_{12}=-J_{12}/|J|$, $|J|=J_{11}J_{22}-J_{12}^{2}$. We have also introduced the following short-hand notation:
$\kappa_{1}^{2}=\frac{q^{2}z_{1}z_{2}}{\varepsilon\varepsilon_{0}}\frac{z_{1}\gamma_{2}+z_{2}\gamma_{1}}{z_{2}J_{11}\gamma_{2}+z_{1}J_{22}\gamma_{1}}.$ (38)
$\mathbf{E}_{2}^{(1)}(0)=-\int\frac{d\mathbb{k}}{(2\pi)^{3}}i\mathbb{k}\psi_{2}^{(1)}{{}_{\mathbb{k}}}.$ (40)
$\mathbf{E}_{2}^{(1)}(0)=-\beta_{2}\mathbb{E},$ (41)
where
$\beta_{2}=\frac{q^{4}z_{1}^{2}z_{2}}{3(\varepsilon\varepsilon_{0})^{2}|J|}\frac{\gamma_{2}(z_{1}J_{22}+z_{2}J_{12})+\gamma_{1}(z_{2}J_{11}+z_{1}J_{12})}{c_{1}\gamma_{2}J_{11}+c_{2}\gamma_{1}J_{22}}\int\frac{d\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{(\mathbb{k}^{2}+\kappa^{2}|\varrho_{\mathbb{k}}|^{2})(\mathbb{k}^{2}+\kappa_{1}^{2}|\varrho_{\mathbb{k}}|^{2})}$ (42)
and
$\kappa^{2}=\frac{q^{2}}{\varepsilon\varepsilon_{0}|J|}(z_{1}^{2}J_{22}+z_{2}^{2}J_{11}+2z_{1}z_{2}J_{12}).$ (43)
To obtain $\beta_{1}$, we have to swap indices 1 and 2 in eq. (42).
The total electric field acting on ion $2$ is
$\mathbb{E}+\mathbf{E}_{2}^{(1)}(0)=(1-\beta_{2})\mathbb{E}.$ (44)
Therefore, the mobility of the ion $a$ ($a=1,2$) is
$m_{a}=\frac{1-\beta_{a}}{\gamma_{a}}.$ (45)
The relaxation contribution to ion mobility is
$m_{a}^{(r)}=-\frac{\beta_{a}}{\gamma_{a}}.$ (46)
Thus, the relaxation contribution to conductivity is determined by the following cumbersome expression
$\Delta\sigma_{r}=\sum\limits_{a}q_{a}^{2}c_{a}m_{a}^{(r)}=-\frac{(z_{1}z_{2})^{3}q^{4}\kappa^{2}}{18\pi\eta\varepsilon\varepsilon_{0}R}\frac{z_{2}\gamma_{1}J_{11}+z_{1}\gamma_{2}J_{22}+(z_{1}\gamma_{1}+z_{2}\gamma_{2})J_{12}}{(z_{1}^{2}J_{22}+z_{2}^{2}J_{11}+2z_{1}z_{2}J_{12})(z_{2}\gamma_{2}J_{11}+z_{1}\gamma_{1}J_{22})}\\ \times\int\frac{d\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{(\mathbb{k}^{2}+\kappa^{2}|\varrho_{\mathbb{k}}|^{2})(\mathbb{k}^{2}+\kappa_{1}^{2}|\varrho_{\mathbb{k}}|^{2})},$ (47)
where we have introduced the ’reduced’ Stokes radius $R=R_{1}R_{2}/(R_{1}+R_{2})$ and used the electroneutrality condition $z_{1}c_{1}=z_{2}c_{2}$ to exclude bulk concentrations of ions in the final expression.

### v2

# Modified Debye-Hückel-Onsager Theory for Electrical Conductivity in Aqueous Electrolyte Solutions. Account of Ionic Charge Nonlocality

@@ Abstract
The paper presents a mean field theory of electrolyte solutions, extending the classical Debye-Hückel-Onsager theory to provide a detailed description of the electrical conductivity in strong electrolyte solutions. The theory systematically incorporates the effects of ion specificity, such as steric interactions, hydration of ions, and their spatial charge distributions, into the mean-field framework. This allows for calculation of ion mobility and electrical conductivity, while accounting for relaxation and hydrodynamic phenomena. At low concentrations, the model reproduces the well-known Kohlrausch’s limiting law. Using the exponential (Slater-type) charge distribution function for solvated ions, we demonstrate that experimental data on the electrical conductivity of aqueous 1:1, 2:1, and 3:1 electrolyte solutions can be approximated over a broad concentration range by adjusting a single free parameter representing the spatial scale of the nonlocal ion charge distribution. Using the fitted value of this parameter at 298.15 K, we obtain good agreement with available experimental data when calculating the electrical conductivity across different temperatures. We also analyze the effects of temperature and electrolyte concentration on the relaxation and electrophoretic contributions to total electrical conductivity, explaining the underlying physical mechanisms responsible for the observed behavior.
@@ I Introduction
Electrolyte solutions [1] are essential for maintaining proper function and balance within living organisms, as they contain ions crucial for cellular processes such as nerve signaling [2, 3], muscle contraction [4, 5], and fluid balance [6]. These solutions can be found in natural sources like fruits and vegetables, as well as in sports drinks and oral rehydration solutions. Beyond biology and the food industry, electrolytes are critical components in technologies with significant societal and environmental impacts. In battery technology, electrolytes facilitate the movement of ions between electrodes, allowing energy storage and release [7, 8, 9, 10, 11]. With growing global demand for fresh water, the role of electrolytes in desalination processes [12, 13] is becoming increasingly important. Electrolytes play a key role in technologies such as reverse osmosis and electrodialysis, facilitating the removal of contaminants from water and the production of clean drinking water. Thus, predicting electrical conductivity in solutions containing strong electrolytes poses significant challenges in physical chemistry and chemical physics, with implications for modern biology, medicine, food industry, agriculture, electronics, water purification, and environmental monitoring technologies.
@@
Debye, Hückel, Onsager, Fuoss, and Falkenhagen [14, 15, 16, 17, 18, 19, 20] laid the foundation for understanding the microscopic theory of electrical conductivity in electrolytes. More recently, renewed interest and progress in this field have arisen from advancements in the statistical theory of liquids and liquid-phase solutions since the publication of these seminal papers. Approaches based on integral equations for distribution functions [21, 22, 23], stochastic density functional theory (SDFT) [24, 25, 26, 21], and self-consistent field theory based on the Poisson-Boltzmann equation [27, 28] have been developed and applied with varying degrees of success to model the electrical conductivity of electrolyte solutions across a broad range of concentrations and temperatures.
@@
Let us review the current state of the art in this field. In paper [29], the author proposed a modified Debye-Hückel-Onsager (DHO) theory, using the concept of smaller-ion shells, which was initially developed by the same author in [30], as a way to enhance the accuracy of predicting activity coefficients for ions. The author’s estimation of the electrical conductivity and activity coefficients of numerous aqueous electrolyte solutions with varying valences, up to concentrations of 0.5-1 mol/l, was achieved without the use of any adjustable parameters and is considered accurate. The paper [25] examines the conductivity of ionic solutions using SDFT combined with a modified Coulomb interaction that takes into account the hard-core repulsion between ions. The authors established that the modified potential suppresses short-range electrostatic interactions present in the original Debye-Hückel-Onsager theory, leading to improved conductivity results. The authors also claim that the conductivity values obtained from the model agree well with the experimental data up to a concentration of 3 mol/l. In a recent paper [26], the same authors extended their study to include multivalent ions and high electric fields, where a deviation from Ohm’s law, known as the Wien effect, occurs. In article [21], the authors explore recent approximations for computing the conductivity and self-diffusion coefficient of electrolytes, with a focus on the specifics of truncating Coulomb interactions at short distances. They compare these new approaches to conventional methods, such as the mean-spherical approximation and mode-coupling theory. The authors also discuss the importance of considering hydrodynamic effects in SDFT and demonstrate how the choice of modified Coulomb interactions can affect the properties of electrolytes. In general, the authors highlight the need for continued development and refinement of SDFT approaches for studying electrolytes and emphasize the valuable insights that can be gained by comparing these methods with other theoretical frameworks. In study [22], the authors developed a new electrical conductivity model for non-associated electrolyte solutions based on the Debye-Hückel-Onsager theory. This model assumes a single cation and a single anion within a continuum medium of the solvent(s), utilizing their crystallographic ionic radii. The authors compared the predictions of their model with experimental measurements of various binary aqueous solutions over a wide temperature range, from 273.15 to 373.15 K, and demonstrated good agreement with the experimental data. In paper [27], the authors provided a comprehensive examination of the unique behavior of electrolyte solutions conductivity when confined to micro- and nanochannels, as compared to bulk phases. They focuses on the reduction of ion mobility with increasing salt concentration, which impacts the conductivity of monovalent solutions as they transition from bulk to confinement in a narrow slit. The authors systematically treated electrophoresis of ions, deriving equations for zeta potentials and ion mobilities, and subsequently derived an elegant analytical expression for bulk conductivity. In the paper [28], the same authors demonstrated that the derived expression accurately describes the conductivity of univalent salt solutions. Note that this model takes into account only a mean-field approximation of electrostatic interactions and uses an adapted classical approach to calculate colloid electrophoresis, as described in [31] and [32]. It is interesting to note that the authors reached an excellent agreement with the experiment without considering the relaxation effect of the salt ions.
@@
Despite several attempts to theoretically describe electrical conductivity in aqueous solutions of strong electrolytes, some important issues remain unclear from the perspective of chemical physics and physical chemistry. Firstly, to the best of our knowledge, the nonlocal nature of the ion charge distribution has not been systematically explored in relation to the electrical conductivity of electrolyte solutions. More specifically, according to quantum theory, the charge of solvated ions (with surrounding solvent molecules in solvation shell) is distributed over a region rather than being concentrated at a single point or spread over a thin shell. In other words, from the perspective of electrostatics, it should behave more like a smooth cloud than a point-like charge or a uniformly charged spherical shell. Based on general considerations, this effect is expected to become significant at sufficiently high solution concentrations, where the interactions between ions can no longer be accurately described as interactions between the point-like charges. It should be noted that recent studies have effectively attempted to account for charge nonlocality by modifying the Coulomb potential at short distances [26, 25, 21]. However, these modifications have not yet been explicitly related to the nature of the charge distribution nonlocality. Note that renormalization of the Coulomb potential at short distances has been systematically considered in plasma physics taking into account quantum effects related to the Heisenberg uncertainty relation [33]. Quantum mechanical calculations of the effective electrostatic interactions between ions in electrolyte solutions present a greater challenge than those in plasma physics, primary due to the presence of multiple electrons in solvated ions. To address this task one can resort to computationally expensive ab initio molecular dynamics simulations. However, for practical calculations, it is also possible to use phenomenological models of the charge distribution of solvated ions adopted in different phenomenological theories [34, 35, 36, 37] or within the microscopic statistical field theory [38, 39, 40, 41, 42]. These approaches lead to a truncation of the Coulomb potential at short distances. Secondly, the effect of ion hydration on the phenomenon of ion transport in a solution under an external electric field is not fully understood. This effect should be analyzed by considering the effective size of the ions (ion hydration radii), taking into account the steric interactions between the ions and their effect on the dielectric constant of the solution (dielectric decrements [43, 44, 45]). For this purpose, steric interactions between ions at close distances could be modeled as interactions of hard spheres with diameter equal to twice the hydration (solvation) radius. It is important to note that existing theoretical models [21, 26, 25] take into account steric effects using crystallographic radii of ions, rather than solvation radii.
@@
This study proposes a modified version of the Debye-Hückel-Onsager (DHO) theory that takes into account ion solvation, steric interactions between ions, and the non-local distribution of charge on solvated ions. We utilize this theory to derive analytical expressions for the electrophoretic and relaxation contributions to ion mobility, as well as to describe the electrical conductivity of electrolyte solutions containing ions of different valencies across a wide range of concentrations and temperatures.
@@ II Statement of problem
For a very diluted solution, the mobility of individual ions approaches a constant limit given by the equation:
@@
For spherical ions, the viscous friction coefficient can be expressed as $\gamma_{a}=6\pi\eta R_{a}$, where $R_{a}$ represents the hydrodynamic (Stokes) radius of the $a$-th ion, and $\eta$ denotes the shear viscosity of the solution. In solutions at finite concentrations each ion is surrounded by a screening charge cloud (atmosphere), which alters its mobility due to two distinct effects. The motion of ions in an external field causes a distortion in the charge distribution within the cloud, leading to an additional ’relaxation’ field acting on the ion. Furthermore, the movement of the ionic atmosphere induces motion in the solvent, resulting in the ion being "carried off". The impact of the first type of contribution on mobility is described as relaxation[46, 47, 20], while the second type is referred to as electrophoretic [46, 47, 27]. In the following, we will present an approach that extends the original Debye-Hückel-Onsager theory [20] to calculate both contributions to the electrical conductivity of strong electrolyte solutions.
@@ III Relaxation contribution
Let us begin by calculating the relaxation contribution to the ion mobilities. We can define the two-point ion concentrations, denoted as $c_{ab}(\mathbb{r}_{a},\mathbb{r}_{b})$. The value $c_{ab}(\mathbb{r}_{a},\mathbb{r}_{b})\text{d}\mathbb{r}_{a}$ represents the number of ions of type $a$ near point $\mathbb{r}_{a}$ when a test ion of type $b$ is fixed at point $\mathbb{r}_{b}$. It is important to note that the ions can be of the same or different types, so we have the following symmetry relation:
@@
Modified Onsager-Fuoss equation. In the configuration space of two ions, $a$ and $b$, in a steady state, which we will only consider in this work, the following continuity equation holds [46, 19, 20]
$\nabla_{a}\cdot\mathbb{j}_{a}+\nabla_{b}\cdot\mathbb{j}_{b}=0,$ (5)
@@
$\bar{\psi}_{b}(\mathbb{r}_{a},\mathbb{r}_{b})=\int\text{d}\mathbb{r}^{\prime}\varrho(\mathbb{r}_{a}-\mathbb{r}^{\prime})\psi_{b}(\mathbb{r}^{\prime},\mathbb{r}_{b})=\varrho\psi_{b}(\mathbb{r}_{a},\mathbb{r}_{b})$ (8)
is the smeared electrostatic potential at point $\mathbb{r}_{a}$ when an ion with total charge $q_{b}$ is located at point $\mathbb{r}_{b}$. The first term in equations (6) and (7) describes the diffusion of ions $a$ even in the absence of an external field. The second term describes the particle flux density due to the action of electrostatic forces from the external field and the ’molecular’ electric field $-\nabla_{a}\bar{\psi}_{b}$ created at point $\mathbf{r}_{a}$ by the distorted ionic cloud, assuming that ion $b$ is located at point $\mathbf{r}_{b}$. The local electrostatic potential, $\psi_{b}=\psi_{b}(\mathbb{r}_{a},\mathbb{r}_{b})$, satisfies the standard Poisson equation
@@
$\rho_{s}(\mathbb{r}_{a},\mathbb{r}_{b})=q_{s}\varrho\bar{c}_{sb}(\mathbb{r}_{a},\mathbb{r}_{b})=q_{s}\int\text{d}\mathbb{r}^{\prime}\varrho(\mathbb{r}_{a}-\mathbb{r}^{\prime})\bar{c}_{sb}(\mathbb{r}^{\prime},\mathbb{r}_{b}).$ (10)
@@
$\varrho f(\mathbb{r})=\int\text{d}\mathbb{r}^{\prime}\varrho(\mathbb{r}-\mathbb{r}^{\prime})f(\mathbb{r}^{\prime}).$ (11)
@@
where $J_{al}=\partial\mu_{a}/\partial c_{l}$, with $\mu_{a}=\mu_{a}(\{c_{l}\})$ being the equilibrium bulk chemical potentials of species which will be specified below. Note also that due to translation invariance all functions should depend only on the vector $\mathbb{r}=\mathbb{r}_{a}-\mathbb{r}_{b}$, so that
@@
The system of coupled equations (15) and (16) form the basis of the modified DHO theory with account for the nonlocal ionic charge distribution, which we will use further to calculate the relaxation contribution to ion mobilities and electrical conductivity.
Onsager-Fuoss approach. In accordance with the Onsager-Fuoss approach[20, 46], considering only the case of very small external electric field, we find the solutions in the form
@@
where we introduced the reciprocal matrix, $\hat{J}_{ab}=(J^{-1})_{ab}$. Therefore, we will be working within the framework of the linear electrostatic approximation, as well as the original DHO theory. The equilibrium electrostatic potential can be obtained from equation
@@
$\nabla^{2}G(\mathbb{r})-\mathcal{K}G(\mathbb{r})=-\frac{1}{\varepsilon\varepsilon_{0}}\varrho(\mathbb{r}).$ (26)
@@
For point-like charges for which $\varrho(\mathbb{r})=\delta(\mathbb{r})$ (or $\varrho_{\mathbb{k}}=1$), we obtain
@@
where $r=|\mathbb{r}|$. Thus, $\kappa^{-1}$ in this case can be interpreted as the screening length [51]. Its dependence on concentrations (and temperature) is determined by the dependence of the bulk chemical potentials on concentrations (and temperature), i.e. $\mu_{a}=\mu_{a}(\{c_{l}\})$ which we will specify below. The equilibrium perturbation concentrations, thereby, are
@@
The steady-state perturbations $\bar{c}_{ab}^{(1)}(\mathbb{r})$ and $\psi_{b}^{(1)}(\mathbb{r})$ which are the scalar values have the structure $(\mathbb{E}\cdot\mathbb{r})f(r)$, where $f(r)$ is the function only of the absolute value of $\mathbb{r}$. Therefore, $\bar{c}_{ab}^{(1)}(\mathbb{r})$ and $\psi_{b}^{(1)}(\mathbb{r})$ are odd functions. Bearing in mind condition (4), we arrive at the following property
@@
$\nabla^{2}\psi_{b}^{(1)}(\mathbb{r})=-\frac{1}{\varepsilon\varepsilon_{0}}\sum\limits_{l}q_{l}\varrho\bar{c}_{lb}^{(1)}(\mathbb{r}).$ (33)
Application to one-component electrolyte solutions. As an application of the proposed theory, let us consider only the case of one-component electrolyte solutions for which indices $a,b=1,2$, so that $q_{1}=z_{1}q$, $q_{2}=-z_{2}q$, $z_{a}=|q_{a}|/q$ is the charge number of ion $a$, $q=e$ is the elementary charge. In this case, only one variable $\bar{c}_{12}^{(1)}(\mathbb{r})\neq 0$. Thus, solving eqs. (32) and (33) by Fourier-method (for derivation details, see Appendix A), we arrive at the following general expression for relaxation contribution to conductivity
$\Delta\sigma_{r}=-\frac{(z_{1}z_{2})^{3}q^{4}\kappa^{2}}{18\pi\eta\varepsilon\varepsilon_{0}R}\frac{z_{2}R_{1}J_{11}+z_{1}R_{2}J_{22}+(z_{1}R_{1}+z_{2}R_{2})J_{12}}{(z_{1}^{2}J_{22}+z_{2}^{2}J_{11}+2z_{1}z_{2}J_{12})(z_{2}R_{2}J_{11}+z_{1}R_{1}J_{22})}\\ \times\int\frac{\text{d}\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{(\mathbb{k}^{2}+\kappa^{2}|\varrho_{\mathbb{k}}|^{2})(\mathbb{k}^{2}+\kappa_{1}^{2}|\varrho_{\mathbb{k}}|^{2})},$ (34)
where, as it follows from eq. (24),
$\kappa^{2}=\frac{q^{2}}{\varepsilon\varepsilon_{0}|J|}(z_{1}^{2}J_{22}+z_{2}^{2}J_{11}+2z_{1}z_{2}J_{12}),$ (35)
and we took into account that $\hat{J}_{11}=J_{22}/|J|$, $\hat{J}_{22}=J_{11}/|J|$, $\hat{J}_{12}=-J_{12}/|J|$, $|J|=J_{11}J_{22}-J_{12}^{2}$. We introduced the following short-hand notation
$\kappa_{1}^{2}=\frac{q^{2}z_{1}z_{2}}{\varepsilon\varepsilon_{0}}\frac{z_{1}R_{2}+z_{2}R_{1}}{z_{2}J_{11}R_{2}+z_{1}J_{22}R_{1}}$ (36)
and the ’reduced’ Stokes radius $R=R_{1}R_{2}/(R_{1}+R_{2})$, and also used the electroneutrality condition $z_{1}c_{1}=z_{2}c_{2}$ to exclude bulk concentrations of ions in the final expression. Eq. (34) is one of the main results of this paper. To apply it to the calculation of the relaxation contribution to conductivity, we need to specify the charge form factor $\varrho_{\mathbf{k}}$ and the matrix elements, $J_{ab}=\partial\mu_{a}/\partial c_{b}$. These quantities will be defined below.
It is instructive to consider the expression for the relaxation contribution, which follows from equation (34), for the case when ions are modeled as point charges, i.e., when $\varrho_{\mathbb{k}}=1$. By taking the integral, we can arrive at
$\Delta\sigma_{r}=-\frac{q^{4}\kappa\,(z_{1}z_{2})^{3}}{72\pi^{2}\eta\varepsilon\varepsilon_{0}R\left(1+\frac{\kappa_{1}}{\kappa}\right)}\frac{z_{2}R_{1}J_{11}+z_{1}R_{2}J_{22}+J_{12}(z_{2}R_{2}+z_{1}R_{1})}{(z_{1}^{2}J_{22}+z_{2}^{2}J_{11}+2z_{1}z_{2}J_{12})(z_{1}R_{1}J_{22}+z_{2}R_{2}J_{11})}.$ (37)